In 5=2x I need to match the expression or equation with an equivalent expression or equation, my choices are:
a.) 1
b.) x
c.) x^5=27
d.) log 2 =5
e.) log 2 5=x
f.)logx5=-2
g.) 2
h.) 0
Thank you

Answer by stanbon(75887)

In 5=2^x I need to match the expression or equation with an equivalent expression or equation, my choices are:
Answer: x = log2(5)
which is "e".
===============
Comment:
Corresponding to every log or ln equation there is
an exponential equation.
Example:
Log Equation: x = log(5)
x is the exponent
10 is the base
5 is the result or anti-log
---
The corresponding exponential equation is:
10^x = 5
Notice that 10 is the base, x is the exponent, and the result is 5.
----------------------------------
Similarly
x = log2(5)
2 is the base
x is the log or exponent
5 is the result or anti-log
---
The corresponding exponential equation is:
2^x = 5
2 is the base
x is the log or exponent
5 is the result or anti-log
If x = log(5
=========================
Cheers,
Stan H.
===============
a.) 1
b.) x
c.) x^5=27
d.) log 2 =5
e.) log 2 5=x
f.)logx5=-2
g.) 2
h.) 0